# Fermi-liquid transport beyond the upper critical field in superconducting La2PrNi2O7 thin films

### v1

@@ Results
Comparing to the previously reported values for La3Ni2O7 and LPNO thin films [13, 15], our extracted $\mu_{0}H_{\rm c2,\perp}(0)$ is significantly lower. We note that in previous reports, a constant resistivity value at a temperature slightly above $T_{\rm c}$ (e.g. 50 K) is used to define the 90% $\rho_{\rm n}$ criterion for $\rho_{\rm xx}(H,T)$ measured at $H\ll H_{\rm c2}$. In LPNO, since the normal-state $\rho_{\rm xx}(T)$ follows a $T^{2}$ behaviour and the $T_{\rm c}\approx$ 40 K is relatively high, the 50-K resistivity value considerably overestimates the true normal-state resistivity at $T\lesssim T_{\rm c}$, thus leading to an overestimation of $H_{\rm c2}(T\lesssim T_{\rm c})$ and $H_{\rm c2}(0)$. It is also likely that there exists a sample variations in $H_{\rm c2}$ (see Supplementary Materials Fig. S1), whose underlying cause requires further investigation. Our current experiment on a LPNO film with an $H_{\rm c2,\perp}$ directly accessible within 53 T provides an accurate quantification of its $\mu_{0}H_{\rm c2,\perp}(0)$, and allows a crude estimate of the superconducting gap magnitude $\Delta_{0}\simeq 9.9$ meV via $\xi_{ab}(0)=\hbar v_{\rm F}/(\pi\Delta_{0})$. We caution here that this $\Delta_{0}$ value should be interpreted as an order-of-magnitude estimate as $v_{\rm F}$ used here is obtained using estimates of $k_{\rm F}$ and $m^{*}$, both of which are associated a sizeable uncertainty, and the level of disorder is assumed to play only a minor role in the extraction of $T_{\rm c}$ and $H_{\rm c2}$ via resistivity measurements.
@@
Figure 3: Upper critical fields of La2PrNi2O7 thin film. $\mu_{0}H_{\rm c2}$ are extracted using the criterion of $\rho_{\rm xx}(\mu_{0}H_{\rm c2})/\rho_{\rm xx}$(53 T) = 0.9. Filled points correspond to the configuration that the magnetic field is applied along the film surface normal (i.e. $\mathbf{H}\parallel c$) and open points the configuration that field applied parallel to the film surface (i.e. $\mathbf{H}\parallel ab$). Dash and dotted lines are fits made using the linearised Ginzburg-Landau formulae (see main text). Inset shows $\mu_{0}H_{c2}$ measured at 35 K with the field oriented close to the in-plane configuration (i.e. $\theta\approx 90^{\circ}$). Solid line is a fit made using the two-dimensional Tinkham model[30]: $\left(\frac{H_{\rm c2}\sin\theta}{H_{\rm c2,\parallel}}\right)^{2}+\left|\frac{H_{\rm c2}\cos\theta}{H_{\rm c2,\perp}}\right|=1$.
@@
It is widely known that the prefactor of $T^{2}$-resistivity and electronic specific heat $\gamma_{0}$ for a variety of correlated electron systems exhibits an empirical relationship[31, 32]:
$A_{2}/\gamma_{0}^{2}\sim 10^{5}\leavevmode\nobreak\ \mu\Omega\leavevmode\nobreak\ \rm{cm\leavevmode\nobreak\ mJ^{-2}\leavevmode\nobreak\ mol^{-2}\leavevmode\nobreak\ K^{-6}}$, known as the Kadowaki-Woods ratio. The high-pressure or epitaxial strain condition required for superconductivity in the bilayer nickelates currently precludes the extraction of $\gamma_{0}$ via direct calorimetry measurements. Alternatively, by assuming the empirical Kadowaki-Woods ratio of $10^{5}\leavevmode\nobreak\ \mu\Omega\leavevmode\nobreak\ \textrm{cm}\leavevmode\nobreak\ \textrm{mJ}^{-2}\leavevmode\nobreak\ \textrm{mol}^{-2}\leavevmode\nobreak\ \textrm{K}^{-6}$ and using the measured $A_{2}=(8.1\pm 1.2)\leavevmode\nobreak\ \textrm{n}\Omega\leavevmode\nobreak\ \textrm{cm}\leavevmode\nobreak\ \textrm{K}^{-2}$, we infer a $\gamma_{0}=(28\pm 4)\leavevmode\nobreak\ \textrm{mJ}\leavevmode\nobreak\ \textrm{mol}^{-1}\leavevmode\nobreak\ \textrm{K}^{-2}$ for the LPNO film studied in this work. We further estimate the quasiparticle effective mass using[33]:
$\gamma_{0}=\gamma\prime\sum_{i}m_{i}^{*}/m_{e},$ (4)
[…]
@@ Discussion
It has been noted that, for a wide variety of unconventional superconductors, the ratio between $T_{\rm c}$ and $T_{\rm F}^{*}$ share a very similar magnitude i.e. $T_{\rm c}/T_{\rm F}^{*}\sim 0.01$ refs. [21, 24, 22, 23]). A survey of $T_{\rm c}$ versus $T_{\rm F}^{*}$ for notable superconductors, including high-$T_{\rm c}$ cuprates[21], iron-based superconductors[22], heavy fermion materials[23], and carbon-based superconductors[24], is shown in Fig. 4. We find that for LPNO the empirical relationship $T_{\rm c}/T_{\rm F}^{*}\sim 0.01$ also holds, suggesting a strongly correlated nature of superconductivity in the bilayer RP nickelates and the magnitude of $T_{\rm c}$ is governed by a universal underlying principle independent of material-specific details.
Figure 4: Superconducting critical temperature $T_{\rm c}$ versus effective Fermi temperature $T_{\rm F}^{*}$ for strongly correlated superconductors. The dotted, dotted-dash, and dashed lines indicate $T_{\rm c}/T_{\rm F}^{*}$ = 0.01, 0.05, and 0.1, respectively. $T_{\rm F}^{*}$ values are extracted assuming a quadratic energy dispersion $E_{\rm F}=\hbar^{2}k_{\rm F}^{2}/(2m^{*})$ and using experimental data of effective mass and carrier density inferred from specific heat, Hall effect, and quantum oscillation measurements (refs. [21, 24, 22, 23] and references therein) or from penetration depth measurements using $E_{\rm F}=(\hbar^{2}/2)(3\pi^{2})^{2/3}n_{\rm s}^{2/3}/m^{*}$ (ref. [38]). The series of points in gradient blue illustrates the effect of hole doping on $T_{\rm c}$ and $T_{\rm F}^{*}$ of La2-xSrxCuO4 (LSCO) with the darker colour indicating a higher doping level up to $x=0.21$ ($T_{\rm c}=27$ K). Error bars for the cuprate materials indicate the variations in $T_{\rm c}$ and $T_{\rm F}^{*}$ with carrier dopings, and for LPNO the variations in $T^{*}_{\rm F}$ inferred from 0-T resistivity measured on multiple samples[15]. MATBG: magic-angle twisted bilayer graphene; TMTSF: tetramethyltetraselenafulvalene; BEDT: bisethylenedithiol; Ba122: BaFe2(As1-xPx)2; STO: SrTiO3; YBCO: YBa2Cu3O6+x; Tl2201: Tl2Ba2CuO6+δ; Tl2223: Tl2Ba2Ca2Cu3O10+δ; Bi2223: Bi2Sr2Ca2Cu3O10+δ; Hg2223: Hg2Ba2Ca2Cu3O10+δ;
@@
Another possibility to account for the contrasting $\rho(T)$ is the qualitative different effect between hydrostatic pressure and epitaxial strain, leading to an opposite change in $c$-axis length and possibly critical distinction in the underlying electronic structure. It has been demonstrated that in epitaxial thin film of La2-xSrxCuO4 grown on LaSrAlO4 near optimal doping ($p=0.15$), $\rho(T)$ evolves from being $T$-linear above $T_{\rm c}$ to being $T$-superlinear as the film thickness reduces from 90 to 3 nm [43], accompanied by a reduction in residual resistivity ratio (RRR=$\rho(\textrm{300\leavevmode\nobreak\ K})/\rho_{0}$) from 6 to 2. We note that in the case of La3Ni2O7 and LPNO, the bulk crystals showing a $T$-linear resistivity have RRR values of 2.0–2.5 (refs.[1, 20, 2]), whereas our thin films show RRR $\approx 4$ (Fig. 1 and Supplementary Materials, Fig. S1). It is thus unlikely that the deviation from a $T$-linear resistivity in LPNO is caused by an enhanced level of disorder, and that the distinction in $\rho(T)$ functional form between that found in bulk crystals and thin films reflects a difference in hole doping and/or strain anisotropy.
Lastly, we comment on the surprising fact that HTSC emerges out of a Fermi-liquid ground state in LPNO. While a $T^{2}$-resistivity is expected for a conventional metallic state as $T\rightarrow 0$, it is generally not expected that such resistivity functional form will hold at high temperatures; instead, $\rho(T)$ in conventional metals typically exhibits a $T$-linear behaviour at intermediate temperature range due to dominant electron-phonon coupling. In the current case of thin-film LPNO, in no temperature range up to 300 K is a $T$-linear resistivity found, implying that the electron-electron interaction in the bilayer nickelates is so dominant that the electron-phonon coupling becomes essentially negligible. This dominant electron-electron coupling is likely to be relevant for understanding the formation of density wave orders in bilayer nickelates [44, 45, 46]. If LPNO is to be mapped to the overdoped regime of high-$T_{\rm c}$ cuprates, it would then be possible to further optimize the $T_{\rm c}$ of RP bilayer nickelates provided a suitable tuning strategy.
@@ Methods
Thin film growth and characterisation
LPNO thin films of approximately 5-nm thickness with a single-unit-cell capping layer of SrTiO3 were grown on SrLaAlO4(001) substrates using pulsed-layer deposition, as detailed in previous work [15]. Gold electrodes of 40-nm thickness with Hall-bar (or van der Pauw) geometry were deposited prior to ozone annealing in a tube furnace. Following transport characterisation measurements conducted using 4He cryostats, the samples were kept cold during storage at liquid nitrogen temperature and transported to high-field facilities at dry-ice temperature.
@@
Ultrasonically bonded aluminum wires were connected to the measurement probe using 25 $\mu$m-diameter gold wires and room-temperature silver paint (DuPont 4929N). Samples were exposed to ambient conditions for less than two hours between taken out of cryogenic storage and loaded into measurement cryostat. Electrical transport measurements with an AC excitation current of 10 $\mu$A were performed in static magnetic fields up to 13 T using a Physical Properties Measurement System by Quantum Design Inc. and in pulsed magnetic fields up to 53.4 T and 64.2 T at the International MegaGauss Science Laboratory in Institute for Solid State Physics, University of Tokyo and Dresden High Magnetic Field Laboratory (HLD-EMFL), Helmholtz-Zentrum Dresden-Rossendorf, respectively.
@@
The magnetoresistivity data at magnetic fields above which $\rho_{\rm xx}(H)$ shows a positive curvature is fitted to: $\rho(H,T)=\rho(0,T)+\beta H^{2}$ for extracting the zero-field resistivity $\rho_{\rm xx}(0,T)$ below $T_{\rm c}$. We further fine-tune the extracted $\rho_{\rm xx}(0,T)$ values to achieve a scaling collapse of the Kohler plot for $\rho_{\rm xx}(H,T)$ measured over the entire temperature range. The adjustment of $\rho_{\rm xx}(0,T)$ required for Kohler scaling shown in Fig. 2b is less than 0.5 $\mu\Omega\leavevmode\nobreak$cm for all temperatures.
@@
We use the 2D free electron model to extract the normal-state electronic structure parameters. The 2D carrier density ($n_{\rm 2D}$) is calculated using:
$n_{\rm 2D}=n_{\rm H}\times d$ (5)
where $d$ is the distance between Ni-O bilayers in adjacent unit cells. The Fermi wavevector, Fermi velocity, Fermi energy, and effective Fermi temperature are then calculated using:
$k_{\rm F}=(2\pi n_{\rm 2D})^{1/2},$ (6)
$v_{\rm F}=\hbar k_{\rm F}/m^{*},$ (7)
$E_{\rm F}=\hbar^{2}k_{\rm F}^{2}/(2m^{*}),$ (8)
$T_{\rm F}^{*}=\frac{\hbar^{2}\pi n_{\rm 2D}}{m^{*}k_{\rm B}},$ (9)
where $\hbar$ and $k_{\rm B}$ are the reduced Planck and Boltzmann constant, respectively, and $m^{*}/m_{e}$ as estimated using the measured $A_{2}$ coefficient and the empirical Kadowaki-Woods ratio of $10^{5}\leavevmode\nobreak\ \mu\Omega\leavevmode\nobreak\ \textrm{cm}\leavevmode\nobreak\ \textrm{mJ}^{-2}\leavevmode\nobreak\ \textrm{mol}^{-2}\leavevmode\nobreak\ \textrm{K}^{-6}$.
@@
Y.-T.H. and Y.Y. conceived the project. Y.Y., Y.-T.H. and H.Y.H. designed the project. Y.L., Y.T. and Y.Y. synthesised and characterised the bilayer nickelate thin films under H.Y.H.’s supervision. Y.-T.H., Y.K., T.K. and V.S. performed the high-field magnetotransport measurements. Y.-T.H. analysed the data and wrote the manuscript with input from all authors.

### v2

@@ Results
Comparing to the previously reported values for La3Ni2O7 and LPNO thin films [14, 16], our extracted $\mu_{0}H_{\rm c2,\perp}(0)$ is significantly lower. We note that in previous reports, a constant resistivity value at a temperature slightly above $T_{\rm c}$ (e.g. 50 K) is used to define the 90% $\rho_{\rm n}$ criterion for $\rho_{\rm xx}(H,T)$ measured at $H\ll H_{\rm c2}$. In LPNO thin films, since the normal-state $\rho_{\rm xx}(T)$ follows a $T^{2}$ behaviour and the $T_{\rm c}\simeq$ 40 K is relatively high, the 50-K resistivity value considerably overestimates the true normal-state resistivity at $T\lesssim T_{\rm c}$, thus leading to an overestimation of $H_{\rm c2}(T\lesssim T_{\rm c})$ and $H_{\rm c2}(0)$. It is also likely that there exists a sample variations in $H_{\rm c2}$ (see Supplementary Materials Fig. S9), whose underlying cause requires further investigation. Our current experiment on a LPNO film with an $H_{\rm c2,\perp}$ directly accessible within 53 T provides an accurate quantification of its $\mu_{0}H_{\rm c2,\perp}(0)$, and allows a crude estimate of the superconducting gap magnitude $\Delta_{0}\simeq 6.3$ meV via $\xi_{ab}(0)=\hbar v_{\rm F}/(\pi\Delta_{0})$. We caution here that this $\Delta_{0}$ value should be interpreted as an order-of-magnitude estimate as $v_{\rm F}$ used here is obtained using estimates of $E_{\rm F}$ and $\bar{m}^{*}$, both of which are associated a sizeable uncertainty, and the level of disorder is assumed to play only a minor role in the extraction of $T_{\rm c}$ and $H_{\rm c2}$ via resistivity measurements. Nonetheless, we note that recent tunneling experiments [39] have reported $\Delta_{0}=(6-20)$ meV for thin-film LPNO, broadly in agreement with our estimate.
@@
Figure 3: Upper critical fields of La2PrNi2O7 thin film. $\mu_{0}H_{\rm c2}$ are extracted using the criterion of $\rho_{\rm xx}(\mu_{0}H_{\rm c2})/\rho_{\rm xx}$(53 T) = 0.9. Filled points correspond to the configuration that the magnetic field is applied along the film surface normal (i.e. $\mathbf{H}\parallel c$) and open points the configuration that field applied parallel to the film surface (i.e. $\mathbf{H}\parallel ab$). Dash and dotted lines are fits made using the linearised Ginzburg-Landau formulae (see main text). Inset shows $\mu_{0}H_{c2}$ measured at 35 K with the field oriented close to the in-plane configuration (i.e. $\theta\simeq 90^{\circ}$). Solid line is a fit made using the 2D Tinkham model[35]: $\left(\frac{H_{\rm c2}\sin\theta}{H_{\rm c2,\parallel}}\right)^{2}+\left|\frac{H_{\rm c2}\cos\theta}{H_{\rm c2,\perp}}\right|=1$.
@@
It is known that the prefactor of $T^{2}$-resistivity and electronic specific heat $\gamma_{0}$ for a variety of correlated electron systems exhibits an empirical relationship[40, 41]: $A_{2}/\gamma_{0}^{2}\simeq 10\penalty 10000\ \mu\Omega\penalty 10000\ \rm{cm\penalty 10000\ mol^{2}\penalty 10000\ K^{2}\penalty 10000\ J^{-2}}$, known as the Kadowaki-Woods ratio. The high-pressure or epitaxial strain condition required for superconductivity in the bilayer nickelates currently precludes the extraction of $\gamma_{0}$ via direct calorimetry measurements. Alternatively, by assuming the empirical Kadowaki-Woods ratio of $10\penalty 10000\ \mu\Omega\penalty 10000\ \rm{cm\penalty 10000\ mol^{2}\penalty 10000\ K^{2}\penalty 10000\ J^{-2}}$ and using the measured $A_{2}=(8.1\pm 1.2)\penalty 10000\ \textrm{n}\Omega\penalty 10000\ \textrm{cm}\penalty 10000\ \textrm{K}^{-2}$, we infer a $\gamma_{0}=(28\pm 4)\penalty 10000\ \textrm{mJ}\penalty 10000\ \textrm{mol}^{-1}\penalty 10000\ \textrm{K}^{-2}$ for the LPNO film studied in this work. We further estimate the quasiparticle effective mass using[42]:
$\gamma_{0}=\gamma\prime\sum_{i}m_{i}^{*}/m_{\rm e},$ (4)
where $\gamma\prime=\frac{\pi N_{\rm A}k_{\rm B}^{2}m_{\rm e}}{3\hbar^{2}}a^{2}=1.39$ mJ mol-1 K-2 ($a=3.756$ Å is the in-plane lattice constant) and $m^{*}_{i}$ is the effective mass of the $i^{\rm th}$ distinct Fermi pocket in the first Brillouin zone (BZ). Given that there are two Ni-O layers in the unit cell of LPNO, two sheets of the Fermi surface are expected in the BZ (i.e. $i=2$) and verified by photoemission experiment [28]. This yields an average effective mass $\bar{m}^{*}=(10\pm 3)\penalty 10000\ m_{e}$ for thin-film LPNO, comparable to that found in overdoped cuprates [43, 44, 45, 46]. We further estimate an effective Fermi temperature $T_{\rm F}=(2330\pm 590)$ K using the carrier density and Fermi energy inferred from photoemission data[28] on thin-film LPNO (see details of the uncertainty estimations in $\bar{m}^{*}$ and $T_{\rm F}$ in Supplementary Materials Sec. J. and Sec. G). Table 1 summarises the key normal-state and superconducting parameters of thin-film LPNO extracted from our transport experiment.
@@ Discussion
It has been noted that, for a wide variety of unconventional superconductors, the ratio between $T_{\rm c}$ and $T_{\rm F}$ share a very similar magnitude i.e. $T_{\rm c}/T_{\rm F}\sim 0.01$ (refs. [22, 25, 23, 24]). A survey of $T_{\rm c}$ versus $T_{\rm F}$ for notable superconductors, including high-$T_{\rm c}$ cuprates[22], iron-based superconductors[23], heavy fermion materials[24], and carbon-based superconductors[25], is shown in Fig. 4. We find that for LPNO the empirical relationship $T_{\rm c}/T_{\rm F}\sim 0.01$ also holds, suggesting a strongly correlated nature of superconductivity in the bilayer RP nickelates and the magnitude of $T_{\rm c}$ is governed by a universal underlying principle independent of material-specific details.
Figure 4: Superconducting critical temperature $T_{\rm c}$ versus effective Fermi temperature $T_{\rm F}$ for strongly correlated superconductors. The dotted, dotted-dash, and dashed lines indicate $T_{\rm c}/T_{\rm F}$ = 0.01, 0.05, and 0.1, respectively. For the referenced materials, $T_{\rm F}$ values are extracted assuming a quadratic energy dispersion $E_{\rm F}=\hbar^{2}k_{\rm F}^{2}/(2m^{*})$ and using experimental data of effective mass and carrier density inferred from specific heat, Hall effect, and quantum oscillation measurements (refs. [22, 25, 23, 24] and references therein) or from penetration depth measurements using $E_{\rm F}=(\hbar^{2}/2)(3\pi^{2})^{2/3}n_{\rm s}^{2/3}/m^{*}$ (ref. [47]). The series of points in gradient blue illustrates the effect of hole doping on $T_{\rm c}$ and $T_{\rm F}$ of La2-xSrxCuO4 (LSCO) with the darker colour indicating a higher doping level up to $x=0.21$ ($T_{\rm c}=27$ K). Error bars for the cuprate materials indicate the variations in $T_{\rm c}$ and $T_{\rm F}$ with carrier dopings, and for LPNO the variations in $T^{*}_{\rm F}$ inferred from the upper and lower limits of $\bar{m}^{*}$ estimates. MATBG: magic-angle twisted bilayer graphene; TMTSF: tetramethyltetraselenafulvalene; BEDT-TTF: bisethylenedithiol-tetrathiafulvalene; Ba122: BaFe2(As1-xPx)2; STO: SrTiO3; YBCO: YBa2Cu3O6+x; Tl2201: Tl2Ba2CuO6+δ; Tl2223: Tl2Ba2Ca2Cu3O10+δ; Bi2223: Bi2Sr2Ca2Cu3O10+δ; Hg2223: Hg2Ba2Ca2Cu3O10+δ;
Our high-field experiment demonstrates that the low-$T$ normal state of compressively strained LPNO thin film exhibits the defining characteristics of Fermi-liquid transport. The dichotomy of normal-state resistivity between pressurised single crystal and strained thin film of superconducting bilayer nickelates thus raises a fundamental question of the appropriate electronic ground state description therein. One possible cause for the contrasting normal-state transport is the impact of effective carrier doping[17]. It is well established that in unconventional superconductors with a superconducting dome, the $\rho(T)$ functional form depends strongly on the effective doping level; a predominantly $T$-linear resistivity is often found near the optimal doping level, contrasting a $T^{2}$ resistivity in the overdoped region on its phase diagram [26, 48, 30, 49, 50]. In such a scenario, our current finding then suggests a possible mapping of epitaxially strained LPNO to the overdoped regime in its putative superconducting dome. Recent transport experiment on La2PrNi2O7-δ with a controlled oxygen deficiency $\delta$ (ref. [19]), however, shows an absence of $T$-linear resistivity with $0.65\leq\delta\leq 0$, implying that an alternative doping strategy is required to explore the putative strange-metal transport in LPNO thin films.
@@
Lastly, we comment on the surprising fact that high-temperature superconductivity emerges out of a Fermi-liquid ground state in LPNO thin films. While a $T^{2}$-resistivity is expected for a conventional metallic state as $T\rightarrow 0$, it is generally not expected that such resistivity functional form will hold at high temperatures; instead, $\rho(T)$ in conventional metals typically exhibits a $T$-linear behaviour at intermediate temperature range between $T_{\rm D}/3\lesssim T\lesssim T_{\rm D}$ due to dominant electron-phonon coupling (for bulk La3Ni2O7, a Debye temperature $T_{\rm D}=383$ K is reported for polycrystals [52]). In the current case of thin-film LPNO, in no temperature range up to 300 K is a $T$-linear resistivity found, implying that the electron-electron interaction therein is so dominant that the electron-phonon coupling becomes essentially negligible. This dominant electron-electron coupling is likely to be relevant for understanding the formation of density wave orders in bilayer nickelates [53, 54, 55]. To the best of our knowledge, the only other material with an apparent $T^{2}$-resistivity up to room temperature is the electron-doped cuprates $R_{2-x}$Ce2CuO4 ($R=$ La, Nd, Pr), whose origin remains to be understood [56]; compared to thin-film LPNO, however, the normal-state resistivity of electron-doped cuprates shows no sign of saturation up to 700 K (ref. [57]), thus the mechanisms behind the high-temperature $T^{2}$ resistivity in electron-doped cuprates and thin-film LPNO are presumably distinct. If the compressively strained LPNO thin films are to be mapped to the overdoped regime of its putative superconducting dome, it would then be possible to further optimize the $T_{\rm c}$ of RP bilayer nickelates provided a suitable tuning strategy.
@@ Methods
Thin film growth and device fabrication
LPNO thin films of approximately 5-nm thickness with a single-unit-cell capping layer of SrTiO3 were grown on SrLaAlO4(001) substrates using pulsed-layer deposition, as detailed in previous work [16], with the crystal structure of thin films confirmed by x-ray diffraction (Supplementary Materials Fig. S1. Gold electrodes of 40-nm thickness with Hall-bar (or van der Pauw) geometry were deposited prior to ozone annealing in a tube furnace (Supplementary Materials Fig. S2. Following transport characterisation measurements conducted using 4He cryostats, the samples were kept cold during storage at liquid nitrogen temperature and transported to high-field facilities at dry-ice temperature.
@@
[…]
@@
The magnetoresistivity data at magnetic fields above which $\rho_{\rm xx}(H)$ shows a positive curvature is fitted to: $\rho(H,T)=\rho(0,T)+\beta H^{2}$ for extracting the zero-field resistivity $\rho_{\rm xx}(0,T)$ below $T_{\rm c}$. We further adjust the extracted $\rho_{\rm xx}(0,T)$ values to achieve a scaling collapse of the Kohler plot for $\rho_{\rm xx}(H,T)$ measured over the entire temperature range. Details of the normal-state resistivity extractions are described in Supplementary Materials Sec. D. The adjustment of $\rho_{\rm xx}(0,T)$ required for Kohler scaling is typically less than 1.0 $\mu\Omega\penalty 10000\$cm.
@@
Electronic specific heat ($\gamma_{0}$) and average effective mass ($\bar{m}^{*}$) are calculated using the following equations:
$\frac{A_{2}}{\gamma_{0}^{2}}=10\penalty 10000\ \mu\Omega\penalty 10000\ \textrm{cm}\penalty 10000\ \textrm{mol}^{2}\penalty 10000\ \textrm{K}^{2}\penalty 10000\ \textrm{J}^{-2},$ (5)
$\gamma_{0}=2\left(\frac{\pi N_{\rm A}k_{\rm B}^{2}}{3\hbar^{2}}a^{2}\right)\bar{m}^{*},$ (6)
where $N_{\rm A}$ is the Avogadro number, $k_{\rm B}$ the Boltzmann constant, $\hbar$ the reduced Planck constant, and $a$ the in-plane lattice constant [16]. Effective Fermi temperature ($T_{\rm F}$) is estimated using photoemission data via two approaches as detailed in Supplementary Materials Sec. G. In short, the first approach fits the energy dispersion $E(k)$ of the $\alpha$-band near the Fermi level using a parabolic function and estimate the Fermi energy ($E_{\rm F}$) using the difference between the band bottom and Fermi level. $T_{\rm F}$ and Fermi velocity ($v_{\rm F}$) are then calculated using:
$T_{\rm F}=E_{\rm F}/k_{\rm B},$ (7)
$v_{\rm F}=\sqrt{2E_{\rm F}/\bar{m}^{*}}.$ (8)
The second approach calculates the average Fermi wavevector ($\bar{k}_{\rm F}$) using a weighted sum of measured $k_{\rm F}$ for the $\alpha$- and $\beta$-pocket by their carrier density:
$\bar{k}_{\rm F}=\frac{n_{\alpha}k_{\rm F,\alpha}+n_{\beta}k_{\rm F,\beta}}{n_{\alpha}+n_{\beta}},$ (9)
$E_{\rm F}=\frac{\hbar^{2}\bar{k}_{\rm F}^{2}}{2\bar{m}^{*}}.$ (10)
A good agreement is found between the $T_{\rm F}$ values estimated by these two approaches.
@@
Y.T.H. and Y.Y. conceived the project. Y.Y., Y.T.H. and H.Y.H. designed the project. Y.L., Y.T. and Y.Y. synthesised and characterised the bilayer nickelate thin films under H.Y.H.’s supervision. Y.T.H., Y.K., T.K. and V.S. performed the high-field magnetotransport measurements. B.Y.W. and Z.X.S. analysed the photoemission data. Y.T.H. analysed the transport data and wrote the manuscript with input from all authors.